# Choose Wisely: Data-Enabled Predictive Control for Nonlinear Systems Using Online Data Selection

### v1

@@ 3 Data Selection
At the heart of Select-DeePC lies the data selection method, which selects data points (representing trajectories $\tau_{i}$) from a large (offline) data set. Since the subproblem solved by Select-DeePC relies on linearly combining trajectories, the subset of data should exhibit predominantly linear dynamics. The suggested approach in the literature is to use a sliding window over past input-output measurements. We maintain the view that $\tilde{\mathcal{D}}$ should be chosen according to some spatial distance metric (in trajectory space) as opposed to temporal proximity (such as a sliding window) to the current linearization point $\tilde{\tau}$.
In the following, we will present two selection methods, norm-based and manifold-embedding-based selection, while placing emphasis on the modular nature of Select-DeePC. While norm-based selection is efficient to implement and allows the user to easily update the data set online, manifold-embedding-based selection enables the distance computation in a lower dimensional embedding space which lessens the impact of the curse of dimensionality but comes with increased offline computation cost.
3.1 Norm-based selection
Norm-based data selection selects from a data set $\mathcal{D}$ the $N_{\text{cols}}$ closest (in space) trajectories by computing the norm of the differences between the open-loop solution $\tilde{\tau}$ and the trajectories in the data set $\mathcal{D}$. An implementation is given in Algorithm 2
@@ 3.1 Norm-based selection
While this approach is intuitive, it is common knowledge that norm-based distances fall flat in high dimensional spaces due to the curse of dimensionality (Aggarwal et al., 2001). Since we are computing norm-based distances of trajectories, i.e. vectors with dimension $(T_{\text{p}}+T_{\text{f}})(m+p)$, this method is not an exception to that.
3.2 Low Dimensional Representation using Manifold Learning
@@
The field of Manifold Learning (Nonlinear Dimensionality Reduction) concerns itself with the compression of high-dimensional data onto manifolds with lower intrinsic dimension than the original data by using (nonlinear) projection techniques (Lee & Verleysen, 2007; Jia et al., 2022). Manifold Learning techniques rely on the manifold hypothesis (Fefferman et al., 2016), which postulates that the intrinsic dimensionality of a data set $\mathcal{D}$ is much lower than the dimensionality of the individual data points. For instance, if the system under consideration is a controllable and deterministic LTI system, then Willems’ Fundamental Lemma states that the trajectories span a manifold (in particular a subspace) that can at most have $n+T_{\text{f}}\cdot m$ degrees of freedom (2005), which is generally less than the trajectory dimensionality $(T_{\text{p}}+T_{\text{f}})(m+p)$. Here, $n$ denotes the dimension of the processes latent state, $p$ is the dimension of the measurement vector and $m$ indicates the dimension of the input signal. Motivated by this observation, we wish to find a lower dimensional representation of the data where the similarities can be computed without incurring the curse of dimensionality.
@@
Isomap, introduced by Tenenbaum et al. (2000), is an unsupervised manifold learning method which aims to find an embedding of the data in a Euclidean space which preserves geodesic distances of a neighborhood graph of the data constructed in the data space. This comes with increased computational cost compared to norm based data selection but can offer inter data point distances which more accurately reflect the nonlinear nature of the data manifold in trajectory space. Also notice that the bulk of this added cost is a one time upfront cost (computing the embedding), which can be performed offline. For algorithmic details, see Appendix C.
By considering the reconstruction error of the data, an estimate of the dimensionality of the data can be found. This dimensionality can be interpreted as the intrinsic dimensionality of the underlying system. See Appendix C for plots demonstrating the diminishing returns of higher embedding dimensions. The relative contrast of a data set $\Delta=(d_{\text{max}}-d_{\text{min}})/d_{\text{min}}$, with $d_{\text{max/min}}$ indicating the maximum and minimum norm values of the data set respectively, is a helpful surrogate to determine how distinct each data point appears under a given distance metric. Figure 3 shows the relative contrast of a data set collected from the rocket simulator in Section 4 as a function of trajectory dimensionality when embedded using Isomap. It compares relative contrast given a fixed embedding dimensionality as well as a varying one, which increases according to $n+T_{\text{f}}\cdot m$, to the relative contrast of the $L_{1}$ norm. While a varying embedding dimensionality offers greater relative contrast for small trajectory lengths, we encounter diminishing returns for large trajectory lengths.
Figure 3: Relative contrast with Isomap embedding compared to $L_{1}$ selection. Both Isomap methods have better relative contrast compared to $L_{1}$ selection. Furthermore, we observe that increasing the embedding dimension further leads to diminishing returns in relative contrast compared to a fixed embedding dimension.
@@
Algorithm 3 Manifold-Embedding-based Data Selection
@@
The pseudocode for manifold-embedding-based data selection using the Isomap algorithm is outlined in Algorithm 3, the details of embed using Isomap can be found in Appendix C. Notice that the embedding of the data set in line 2 can be cached/computed offline. Online, the algorithm then proceeds to embed the query trajectory $\tilde{\tau}_{i}$ and computes its similarities in the lower dimensional embedding space. Then, similarly to Algorithm 2, the relevant trajectories are returned in the trajectory space, after which the implicit predictor for DeePC is constructed using the retrieved non-embedded data points.
@@ 4 Results and Discussion
Simulation experiments were performed to validate the efficacy of Select-DeePC and to benchmark it against classical DeePC, which utilizes the entire data set, as well as the Time-Windowed DeePC variant proposed by Berberich et al. (2022), which constructs the implicit predictor by continuously updating the data set with new input-output data and discarding the most out-of-date measurement.
@@ 4.1 Landing a Reusable Rocket
We use a planar vertical takeoff, vertical landing rocket to demonstrate the performance of Select-DeePC on an open-loop unstable nonlinear system. We show that Select-DeePC successfully copes with the nonlinear dynamics and manages to stabilize the system at a given setpoint. A custom environment [^1] https://gitlab.ethz.ch/bsaverio/coco-project for the gymnasium suite (Kwiatkowski et al., 2024) was used as a simulation platform. The rocket measurement $y\in\mathbb{R}^{6}$ includes the 2 dimensional position, the corresponding velocities, the heading, and the angular rate of the rocket. The rocket control input $u\in\mathbb{R}^{3}$ is the main engine thrust, the angle of the main engine gimbal, and the thrust magnitude of the side thrusters. For a detailed description of the environment, we refer to (Cummins et al., 2024).
The right column in Figure 1 shows the simulator setup and specifically, in the bottom figure, three closed-loop rocket trajectories. One for Select-DeePC, one for standard DeePC and one for Time-Windowed DeePC respectively. While both the standard DeePC and the Time-Windowed DeePC controllers diverge, Select-DeePC manages to regulate the system to the desired setpoint and stabilize it there.
Ablation Study. We compare the performance of Select-DeePC using the two presented selection methods on two different data sets. Both data sets were collected using input sequences of the form
@@
The data set IID Data uses $(a,b)=(0,1)$ and Random Walk Data uses $(a,b)=(1,0.1)$. To ensure that a large region of the trajectory space was covered by the collected data, 100 simulations were run with a maximum of 100 timesteps each, resulting in data sets with $5000$$ data points. As shown later, the level of nonlinearity in the data has a significant effect on the performance of standard DeePC. Due to the high inertia of the rocket, the IID Data predominantly exhibits behaviors that can be represented reasonably well by a linear model, and hence, standard DeePC can be applied. However, this comes at a performance penality as nonlinear effects are not accounted for. We observe that Select-DeePC still outperforms standard DeePC on IID Data and that data selection can still lead to performance increases if the data set does not fully capture the nonlinear behavior of the underlying system. The random walk structure of the input sequence in Random Walk Data, on the other hand, ensures that the nonlinearities of the system are sufficiently excited. This makes standard DeePC fail but provides Select-DeePC with more global information about the nonlinear system.
The comparisons of Select-DeePC on the two data sets are two-fold. First, the resulting predictive accuracy of the selection methods are compared. Furthermore, a comparison of the realized closed-loop cost is given as a function of selection method and data set. As a benchmark, random sketching of the full data set is used at each decision moment.
Figure 4: Residual between the ground-truth and the least-squares prediction as a function of trajectories used in the predictor. The selection methods compared here are using random sketching of the data set, $L_{1}$-norm based and Isomap-embedding based selection. Notice that data selection leads to lower cumulative residual than using the full data set (right most data point).
The predictive accuracy of the methods was compared by computing the least squares solution of the predictor used in the DeePC subproblem as a function of the number $N_{\text{cols}}$ of trajectories in the selected set of data. For each trajectory in the holdout data set, constructed from trajectory-segments of a closed-loop trajectory, the $N_{\text{cols}}$ closest trajectories (in spatial $L_{1}$-norm or using Isomap-embedding) under the given selection method were selected and the prediction of future measurements $\hat{y}_{\text{f}}$, given the corresponding future input sequence $u_{\text{f}}$ was computed using Equation 11. The plots in Figure 4 show the cumulative residuals between the predicted future trajectory and the ground truth realization $\|\hat{y}_{\text{f}}-y_{\text{f}}\|_{2}$ over the entire validation set.
@@
Figure 5: Closed-loop cost as a function of data subset cardinality on two different data sets. Instances where the selection methods realized infinite cost (due to instability) have been omitted and the remaining finite cost data points are shown. Selecting data points outperforms using the full data on all data sets.
Figure 5 shows the closed-loop performance of Select-DeePC on the setpoint tracking task depicted in the right column of Figure 1 on both IID Data and Random Walk Data. We again observe on both data sets that selecting data outperforms using the full data set in terms of closed-loop performance, as a lower cost could be achieved. Using IID Data, the reduction in closed-loop cost is 1.9 while using Random Walk Data, DeePC using the full data set failed to stabilize the rocket.
Coincidentally, we observe that random sketching of the data set only results in marginally worse performance compared to Isomap or $L_{1}$ selection. This could be explained by the fact that the system is still relatively close to linear as the attitude of the rocket does not exceed 60°60°60  ° and the high inertia of the system can help in compensating bad actions taken in one timestep—as the randomly selected data are resampled every timestep, it is unlikely that the sampled data poorly represents the system dynamics in multiple consecutive timesteps.
Computational Cost.
Figure 6: Solve times of Select-DeePC as a function of number of Hankel columns decomposed into data selection time and QP solve time. While the QP solve time increases as the number of Hankel columns increases, the data selection time stays constant. It is only affected by the size of the entire data set which stays constant.
Figure 6 shows the computation times of Select-DeePC as a function of number of subselected trajectories. As postulated in Section 2, the dominant factor in the solve times is solving the quadratic program that scales at least quadratically with $N_{\text{cols}}$, while data selection is a constant time operation with respect to $N_{\text{cols}}$.
@@ 4.2 Planar Robotic Manipulator
We use the Reacher (Kwiatkowski et al., 2024) environment, which simulates a planar robotic manipulator, to demonstrate the ability of Select-DeePC to control a highly nonlinear system and successfully track reference setpoints. Furthermore, we show that Select-DeePC is effective in coping with and respecting a posteriori specified output constraints. The kinematic chain of the manipulator consists of two rigid bodies that are connected by revolute joints which are each actuated by a torque, giving it an action space of $u\in\mathbb{R}^{2}$. The measurement vector consists of the sine and cosine of both joint angles respectively, and the end-effector position in $\mathbb{R}^{2}$. Furthermore, we are also able to measure the angular velocity of each link, resulting in an observation space $y\in\mathbb{R}^{8}$. This system is more challenging to control using DeePC (or any linear control method) since the end-effector position is dictated by the composition of two rotations. In this environment, it is not unreasonable to assume that the angles must exceed $90^{\circ}$ in order to reach a target. Furthermore, for a given end-effector position, the corresponding joint configuration is not necessarily unique.
Ablation Study. We again investigate the effect of number of selected trajectories $N_{\text{cols}}$, as well as selection strategy on the resulting closed loop cost. For the Reacher, data were collected using IID inputs. In total, 200 simulations were run with 200 steps each. Here, only IID inputs were collected instead of random walk because the system has lower inertia compared to the rocket and the nonlinearities could be excited more easily.
@@
Figure 7: Comparing the closed-loop cost incurred as a function of subset cardinality. Clearly there is a sweet spot in how many trajectories should be selected for the predictions. Furthermore, due to the highly nonlinear nature of the problem, distance based data selection clearly outperforms random sketching of the full data set.
Figure 7 again shows the incurred closed-loop cost as a function of number of trajectories used in the implicit predictor. Similarly to the rocket, we observe a decrease in predictive accuracy and, as a consequence, an increase in cost as the number of data points used in the predictor increases. While random sketching of the data set showed decent results in the rocket simulation, it fails completely in the reacher simulation as no accurate predictions could be generated, resulting in the controller not making any progress towards the setpoint and hence accumulating high closed-loop cost.
Figure 8: Closed-loop trajectories for a constrained (dark blue) and unconstrained (blue) DeePC subproblem.
Zero-Shot Generalizability of Select-DeePC. Figure 8 shows two closed-loop trajectories of Select-DeePC where the controller was tasked with tracking a set of consecutive end-effector references. We observe that Select-DeePC is able to generalize to different tasks across the entire trajectory manifold of the Reacher by reusing the same data set and without structural changes to the controller. While one simulation was left unconstrained, a constraint of $y<0.1$ was added to the DeePC subproblem of the other one, indicated by the red region. The structure of Select-DeePC effectively decouples the dynamics learning/training process from the controller design. Indeed, both the constrained and unconstrained tasks use the same data set which is unaware of the structure of the predictive control problem. This allows us to easily adapt the cost function or constraint terms for the task at hand in zero-shot fashion without requiring task specific data or additional online learning.
@@ 4.3 Cart-Pole inverted pendulum swing-up
Figure 9: Closed-loop trajectory for an inverted pendulum swing-up. Select-DeePC (blue) is successful in transitioning from the downward facing stable equilibrium to the upward facing unstable equilibrium. Standard DeePC (orange) on the other hand fails to transition between the equilibria. The dashed blue line indicates the open-loop prediction of Select-DeePC at the four snapshot time instances, t=15, 22, 32 and 120.
Finally, we show that Select-DeePC is capable of performing a cart-pole inverted pendulum swing-up, which is a standard benchmark for nonlinear MPC methods. This test demonstrates that Select-DeePC successfully plans trajectories for the nonlinear dynamics using its implicit linearization in trajectory-space. The input $u\in\mathbb{R}$ of the inverted pendulum is a force acting on the cart along the $x$-axis. Furthermore, the measurement $y\in\mathbb{R}^{5}$ consists of the cart position, the sine and cosine of the angle between the pendulum and the vertical axis, as well as linear and angular velocities of the cart and pendulum respectively.
Due to the unstable nature of the desired operating point and the system starting in a stable equilibrium position, data was gathered in closed-loop. Specifically, the data set consists of 200 demonstrations of successful but suboptimal swing-ups, collected using an energy based feedback controller (Chatterjee et al., 2002).
Figure 9 shows the closed-loop trajectories of both Select-DeePC and standard DeePC when tasked with tracking a setpoint at the upward facing, unstable equilibrium while starting from the downward facing stable equilibrium. While standard DeePC is unable to perform the swing-up and oscillates around the stable equilibrium, Select-DeePC is successfully able to learn from the demonstrations in the data set and performs a swing-up to the desired vertical position and subsequently stabilizes the pendulum there. It is worthwhile to point out that the controller does not simply select the top trajectory in the selected data set. Instead, Select-DeePC linearly combines several trajectory segments at each iteration of the solver to produce the swing-up (c.f. Figure 10). This is also evident from the fact that the horizon in the predictive controller is shorter than the episode length, thus the solver never has access to a full swing-up trajectory.
Figure 10: Number of trajectories used to construct the final prediction at each sampling instant of the inverted pendulum swing-up.

### v2

@@ 3 Data Selection
At the heart of Select-DPC lies the data selection method, which selects data points (representing trajectories $\tau_{i}$) from a large (offline) data set. Since the subproblem solved by Select-DPC relies on linearly combining trajectories, the subset of data should exhibit predominantly linear dynamics. The suggested approach in the literature is to use a sliding window over past input-output measurements. We propose that $\tilde{\mathcal{D}}$ should be chosen according to some spatial distance metric (in trajectory space) as opposed to temporal proximity (such as a sliding window) to the current linearization point $\tilde{\tau}$.
In the following, we will present two selection methods, norm-based and manifold-embedding-based selection, while placing emphasis on the modular nature of Select-DPC. While norm-based selection is efficient to implement and allows the user to easily update the data set online, manifold-embedding-based selection enables the distance computation in a lower dimensional embedding space which lessens the impact of the curse of dimensionality but comes with increased offline computation cost.
Norm-based selection
Norm-based data selection selects from a data set $\mathcal{D}$ the $N_{\text{cols}}$ closest (in space) trajectories by computing the norm of the differences between the open-loop solution $\tilde{\tau}$ and the trajectories in the data set $\mathcal{D}$. An implementation is given in Algorithm 2.
@@
While this approach is intuitive, it is common knowledge that norm-based distances fall flat in high dimensional spaces due to the curse of dimensionality Aggarwal et al. (2001). This method is no exception since we are computing norm-based distances of trajectories, i.e. vectors with dimension ${(T_{\text{p}}+T_{\text{f}})(m+p)}$.
Low Dimensional Representation using Manifold Learning
@@
The field of Manifold Learning (Nonlinear Dimensionality Reduction) concerns itself with the compression of high-dimensional data onto manifolds with lower intrinsic dimension than the original data by using (nonlinear) projection techniques Lee & Verleysen (2007); Jia et al. (2022). Manifold Learning techniques rely on the manifold hypothesis Fefferman et al. (2016), which postulates that the intrinsic dimensionality of a data set $\mathcal{D}$ is much lower than the dimensionality of the individual data points. For instance, if the system under consideration is a controllable and deterministic LTI system, then Willems’ Fundamental Lemma states that the trajectories span a manifold (in particular a subspace) that can at most have $n+T_{\text{f}}\cdot m$ degrees of freedom Willems et al. (2005), which is generally less than the trajectory dimensionality $(T_{\text{p}}+T_{\text{f}})(m+p)$. Here, $n$ denotes the dimension of the process’s latent state, $p$ is the dimension of the measurement vector and $m$ indicates the dimension of the input signal. Motivated by this observation, we wish to find a lower dimensional representation of the data where the similarities can be computed without incurring the curse of dimensionality.
@@
Isomap, introduced by Tenenbaum et al. (2000), is an unsupervised manifold learning method which aims to find an embedding of the data in a Euclidean space which preserves geodesic distances of a neighborhood graph of the data constructed in the data space. This comes with increased computational cost compared to norm based data selection but can offer inter data point distances which more accurately reflect the nonlinear nature of the data manifold in trajectory space. Also notice that the bulk of this added cost is a one-time upfront cost (computing the embedding), which can be performed offline. For algorithmic details, see Appendix C.
@@
Algorithm 3 Manifold-Embedding Data Selection
@@
The pseudocode for manifold-embedding-based data selection using the Isomap algorithm is outlined in Algorithm 3, the details of embed using Isomap can be found in Appendix C. Notice that the embedding of the data set in line 2 can be cached/computed offline. Online, the algorithm then proceeds to embed the query trajectory $\tilde{\tau}_{i}$ and computes its similarities in the lower dimensional embedding space. Then, similarly to Algorithm 2, the relevant trajectories are returned in the trajectory space, after which the implicit predictor for DPC is constructed using the retrieved non-embedded data points.
@@ 4 Results and Discussion
Simulation experiments were performed to validate the efficacy of Select-DPC and to benchmark it against classical DeePC Coulson et al. (2019), which utilizes the entire data set, as well as the Time-Windowed DeePC variant proposed by Berberich et al. (2022), which constructs the implicit predictor by continuously updating the data set with new input-output data and discarding the most out-of-date measurement. All experiments were performed on an Apple MacBook Pro with M2 Pro chip and 16GB of RAM. To speed up computation times, an LQ decomposition of the subset of data was performed in the experiments regarding closed-loop cost for large subset sizes Breschi et al. (2023). We use DeePC as the linear DPC solver in Select-DPC.
@@ 4.1 Landing a Reusable Rocket
We use a planar vertical takeoff, vertical landing rocket to demonstrate the performance of Select-DPC on an open-loop unstable nonlinear system. We show that Select-DPC successfully copes with the nonlinear dynamics and manages to stabilize the system at a given setpoint. A custom environment for the gymnasium suite Kwiatkowski et al. (2024) was used as a simulation platform. For a detailed description of the environment, we refer to Cummins et al. (2024).
The right column in Figure 1 shows the simulator setup and specifically, in the bottom figure, three closed-loop rocket trajectories. One for Select-DPC, one for standard DeePC and one for Time-Windowed DeePC respectively. While both the standard DeePC and the Time-Windowed DeePC controllers diverge, Select-DPC manages to regulate the system to the desired setpoint and stabilize it there.
Ablation Study
We compare the performance of Select-DPC using the two presented selection methods on two different data sets. Both data sets were collected using input sequences of the form
@@
The data set IID Data uses $(a,b)=(0,1)$ and Random Walk Data uses $(a,b)=(1,0.1)$. To ensure that a large region of the trajectory space was covered by the collected data, 100 simulations were run with a maximum of 100 timesteps each, resulting in data sets with $5000$$ data points. As shown later, the level of nonlinearity in the data has a significant effect on the performance of standard DeePC. Due to the high inertia of the rocket, the IID Data predominantly exhibits behaviors that can be represented reasonably well by a linear model, and hence, standard DeePC can be applied. However, this comes at a performance penality as nonlinear effects are not accounted for. We observe that Select-DPC still outperforms standard DeePC on IID Data and that data selection can still lead to performance increases if the data set does not fully capture the nonlinear behavior of the underlying system. The random walk structure of the input sequence in Random Walk Data, on the other hand, ensures that the nonlinearities of the system are sufficiently excited. This makes standard DeePC fail but provides Select-DPC with more global information about the nonlinear system.
The comparisons of Select-DPC on the two data sets are two-fold. First, the resulting predictive accuracy of the selection methods are compared. Furthermore, a comparison of the realized closed-loop cost is given as a function of selection method and data set. As a benchmark, random sketching of the full data set at each decision moment is used.
Figure 3: Residual between the ground-truth and the least-squares prediction as a function of trajectories used in the predictor.
The predictive accuracy of the methods was compared by computing the least squares solution of the predictor used in the DPC subproblem as a function of the number $N_{\text{cols}}$ of trajectories in the selected set of data. For each trajectory in the holdout data set, constructed from trajectory-segments of a closed-loop trajectory, the $N_{\text{cols}}$ closest trajectories (in spatial $L_{1}$-norm or using Isomap-embedding) under the given selection method were selected and the prediction of future measurements $\hat{y}_{\text{f}}$, given the corresponding future input sequence $u_{\text{f}}$ was computed using Equation 11. The plots in Figure 3 show the cumulative residuals between the predicted future trajectory and the ground truth realization $\|\hat{y}_{\text{f}}-y_{\text{f}}\|_{2}$ over the entire validation set.
@@
Figure 4: (left) Closed-loop cost as a function of data subset cardinality on two different data sets. Instances where the selection methods realized infinite cost (due to instability) have been omitted and the remaining finite cost data points are shown. (right) Solve times of Select-DPC as a function of number of Hankel columns decomposed into data selection time and QP solve time.
The left plot of Figure 4 shows the closed-loop performance of Select-DPC on the setpoint tracking task depicted in the right column of Figure 1 on both IID Data and Random Walk Data. We again observe on both data sets that selecting data outperforms using the full data set in terms of closed-loop performance, as a lower cost could be achieved. Using IID Data, the reduction in closed-loop cost is 1.9 while using Random Walk Data, standard DeePC using the full data set failed to stabilize the rocket.
Computational Cost
The right plot of Figure 4 shows the computation times of Select-DPC as a function of number of subselected trajectories. As postulated in Section 2, the dominant factor in the solve times is solving the quadratic program that scales at least quadratically with $N_{\text{cols}}$, while data selection is a constant time operation with respect to $N_{\text{cols}}$.
@@ 4.2 Planar Robotic Manipulator
We use the Reacher environment Kwiatkowski et al. (2024), which simulates a planar robotic manipulator, to demonstrate the ability of Select-DPC to control a highly nonlinear system and successfully track reference setpoints. Furthermore, we show that Select-DPC is effective in coping with and respecting a posteriori specified output constraints. The kinematic chain of the manipulator consists of two rigid bodies that are connected by revolute joints which are each actuated by a torque, giving it an action space of $u\in\mathbb{R}^{2}$. The measurement vector consists of the sine and cosine of both joint angles respectively, and the end-effector position in $\mathbb{R}^{2}$. Furthermore, we are also able to measure the angular velocity of each link, resulting in an observation space $y\in\mathbb{R}^{8}$. This system is more challenging to control using DeePC (or any linear control method) since the end-effector position is dictated by the composition of two rotations. In this environment, it is not unreasonable to assume that the angles must exceed $90^{\circ}$ in order to reach a target. Furthermore, for a given end-effector position, the corresponding joint configuration is not necessarily unique.
Figure 5: (left) Closed-loop trajectories for a constrained (dark blue) and unconstrained (blue) DPC subproblem. (right) Comparing the closed-loop cost incurred as a function of subset cardinality. Clearly there is a sweet spot in how many trajectories should be selected for the predictions.
Ablation Study
We again investigate the effect of number of selected trajectories $N_{\text{cols}}$, as well as selection strategy on the resulting closed loop cost. For the Reacher, data were collected using IID inputs. In total, 200 simulations were run with 200 steps each. Here, only IID inputs were collected instead of random walk because the system has lower inertia compared to the rocket and the nonlinearities could be excited more easily.
@@
The right plot of Figure 5 again shows the incurred closed-loop cost as a function of number of trajectories used in the implicit predictor. Similarly to the rocket, we observe a decrease in predictive accuracy and an increase in cost as the number of data points used in the predictor increases. While random sketching of the data set showed decent results in the rocket simulation, it fails completely in the reacher simulation as no accurate predictions could be generated, resulting in the controller not making any progress towards the setpoint and hence accumulating high closed-loop cost.
Zero-Shot Generalizability of Select-DPC
The left plot of Figure 5 shows two closed-loop trajectories of Select-DPC where the controller was tasked with tracking a set of consecutive end-effector references. We observe that Select-DPC is able to generalize to different tasks across the entire trajectory manifold of the Reacher by reusing the same data set and without structural changes to the controller. While one simulation was left unconstrained, a constraint of $y<0.1$ was added to the DPC subproblem of the other one, indicated by the red region. The structure of Select-DPC effectively decouples the dynamics learning/training process from the controller design. Indeed, both the constrained and unconstrained tasks use the same data set which is unaware of the structure of the predictive control problem. This allows us to easily adapt the cost function or constraint terms for the task at hand in zero-shot fashion without requiring task specific data or additional online learning.
@@ 4.3 Cart-Pole inverted pendulum swing-up
Finally, we show that Select-DPC is capable of performing a cart-pole inverted pendulum swing-up, which is a standard benchmark for nonlinear MPC methods. This test demonstrates that Select-DPC successfully plans trajectories for the nonlinear dynamics using its implicit linearization in trajectory-space. The input $u\in\mathbb{R}$ of the inverted pendulum is a force acting on the cart along the $x$-axis. Furthermore, the measurement $y\in\mathbb{R}^{5}$ consists of the cart position, the sine and cosine of the angle between the pendulum and the vertical axis, as well as linear and angular velocities of the cart and pendulum respectively.
Due to the unstable nature of the setpoint and the system starting in a stable equilibrium position, data was gathered in closed-loop. Specifically, the data set consists of 200 demonstrations of successful but suboptimal swing-ups, collected using an energy based feedback controller Chatterjee et al. (2002).
The left plot of Figure 6 shows the closed-loop trajectories of both Select-DPC and standard DeePC when tasked with tracking a setpoint at the upward facing, unstable equilibrium while starting from the downward facing stable equilibrium. While standard DeePC is unable to perform the swing-up and oscillates around the stable equilibrium, Select-DPC is successfully able to learn from the demonstrations in the data set and performs a swing-up to the desired vertical position and subsequently stabilizes the pendulum there. It is worthwhile to point out that the controller does not simply select the top trajectory in the selected data set. Instead, Select-DPC linearly combines several trajectory segments at each iteration of the solver to produce the swing-up (c.f. right plot of Figure 6). This is also evident from the fact that the horizon in the predictive controller is shorter than the episode length, thus the solver never has access to a full swing-up trajectory.
Figure 6: (left) Closed-loop trajectory for an inverted pendulum swing-up. Select-DPC (blue) is successful in transitioning from the downward facing stable equilibrium to the upward facing unstable equilibrium. Standard DeePC (orange) on the other hand fails to transition between the equilibria. The dashed blue line indicates the open-loop prediction of Select-DPC at the four snapshot time instances, t=15, 22, 32 and 120. (right) Number of trajectories used to construct the final prediction at each sampling instant of the inverted pendulum swing-up.